OpenStudy (anonymous):

A certain drug is made from only two ingredients. Compound A and Compound B. There are 5 milliliters of compound A used for every 4 milliliters of compound B. If a chemist wants to make 873 milliliters of the drug, how many milliliters of compound A are needed?

OpenStudy (anonymous):

please help!

OpenStudy (science0229):

From the question, you know that 4A=5B. Also, you know that A+B=873

OpenStudy (science0229):

Do you know how to solve systems of equations?

OpenStudy (anonymous):

I'm not that well at it, so no.

OpenStudy (science0229):

Okay. I'll walk you step by step.

OpenStudy (science0229):

First, for the 2nd equation, A+B=873, do you agree that 4*(A+B)=4*873?

OpenStudy (anonymous):

Yes, I agree

OpenStudy (science0229):

Now, that equation becomes 4A+4B=3492. By the first equation, 4A=5B, the equation changes to 9B=3492. I think you can take it from here.

OpenStudy (anonymous):

Thanks very much!

OpenStudy (science0229):

Welcome :)

OpenStudy (anonymous):

I divide next and then that's my answer, correct?